Given: f(x)=3x-5 and g(x)=-5x+2
1. Find f[g(-1)]
2. Find g[f(4)]

Answer by tutorcecilia(2144)

Given: f(x)=3x-5 and g(x)=-5x+2
1. f[g(-1)]= 3(-5x+2)-5 [Substitute "-5x+2" for x]
f[g(-1)]= 3(-5(-1)+2)-5 [Substitute "-1" for x]
f[g(-1)]= 3(-5(-1)+2)-5=16
.
2. g[f(4)]= -5(-3x-5)+2 [Substitute "3x-5" for x]
-5(-3(4)-5)+2 [Substitute "4" for x]
-5(-3(4)-5)+2=87